# Efficiently learning depth-3 circuits via quantum agnostic boosting

### v2

@@ Abstract
Using quantum agnostic boosting, we obtain a $n^{O(\log\log n\cdot\log(1/\varepsilon))}$-time algorithm for learning $\textsf{poly}(n)$-sized depth-$3$ circuits (consisting of AND, OR, NOT gates) in the uniform PAC model given quantum examples, which is near-polynomial time for constant $\varepsilon$. Classically, obtaining an algorithm with a similar complexity has been an open question in the PAC model and our work answers this given quantum examples.
@@ 1 Introduction
Classically, a well-known idea to PAC learn depth-$3$ circuits is to consider the formulation of agnostic learning. In particular, works of [Fel09, KMV08] showcased how agnostic learning depth-$2$ circuits, in particular DNF formulas, could yield learning algorithms for depth-$3$ circuits. This naturally motivates the need to understand learning DNFs in the quantum agnostic learning model.
@@
Tomography of quantum states (i.e., learning quantum states in the noise-free model) has been well-studied in quantum computing. Quantum agnostic learning got traction only recently with the works of Grewal et al. [GIKL24b] and Chen et al. [CGYZ25] that considered learning stabilizer states. Subsequently, a few works considered stabilizer product states [GIKL24a], high stabilizer-dimension states [CGYZ25] and product states [BBK+25]. In the pursuit of agnostic learning DNFs, we initiate the problem of agnostic learning phase states (which is incomparable to the works mentioned and which we will discuss in more detail below). We now define the model of agnostic learning specialized to phase states. For the concept class ${\mathcal{C}}\subseteq\{c:{\{0,1\}}^{n}\rightarrow{\{0,1\}}\}$, we denote the phase state $|\phi_{c}\rangle$ corresponding to a function $c\in{\mathcal{C}}$ as
@@ 1.1 Main results
Our main result is to make progress on both of the questions highlighted above. For both the tasks of agnostically learning $\operatorname{poly}(n)$-sized DNFs and PAC learning $\operatorname{poly}(n)$-sized depth-$3$ circuits, we give $n^{O(\log\log n\cdot\log(1/\varepsilon))}$ time quantum algorithms (which for constant $\varepsilon$ can be viewed as “near-polynomial” time in the PAC model). Below, we will discuss these results, starting with our main technical contribution, quantum agnostic boosting.
@@
We formally state this theorem in Section 3. To provide some context, classically, Freund and Schapire [Sch90, Fre95, FS99] proposed the boosting algorithm called AdaBoost that efficiently uses a weak learner as a black-box to construct a strong learner in the usual PAC model. The AdaBoost algorithm by Freund and Schapire was one of the first few theoretical boosting algorithms that was simple enough to be extremely useful and successful in practice [SF12]. Similarly, agnostic boosting has been considered by works of [BDLM01, KMV08, Fel09] wherein they show similar results to boost weak agnostic learners to strong ones.
In the quantum setting, there are only a handful of works that have used boosting in the PAC model [BJ95, AM20, IdW20], and ours is the first work that demonstrates how to perform boosting in the harder model of agnostic learning. [^1] We remark that [CTB24] also discusses an agnostic quantum booster but their model assumes that the unknown quantum state is a function state whereas we make no assumption on our input state. Apart from the context of learning Boolean functions, we believe that our quantum boosting algorithm could have utility for learning more general quantum states. Recently, [AD25] utilized boosting on top of a weak agnostic learner for stabilizer states to give tomography protocols of states promised to have structured stabilizer decompositions.
@@
We formally state this theorem for each class in Section 4. In order to compare with classical results and for ease of exposition, let us consider the parameters above to take values of $2^{k},t,s,1/\varepsilon=\operatorname{poly}(n)$ (which in general are considered the most interesting settings of these parameters for learning). Classically, the SOTA algorithm for learning decision trees in the agnostic model (without membership queries) runs in $n^{O(\log n)}$ [KKMS08] time and with membership queries scales as $\operatorname{poly}(n)$ [GKK08b]. [^3] By membership queries we mean that an algorithm can query an unknown $f:{\{0,1\}}^{n}\rightarrow{\{0,1\}}$ on an $x$ of its choice. In the (uniform) agnostic model, by membership query we mean the following: for every distribution $D:{\{0,1\}}^{n+1}\rightarrow[0,1]$ whose marginal on the first $n$ bits is uniform, the learner can query $x$ and obtain a sample from $b\sim D(x,\cdot)$. The same results hold for juntas since $k$-juntas have size-$2^{k}$ decision trees. In contrast, for DNF formulas, the SOTA algorithm in the agnostic model is $n^{O(\log^{2}n)}$ [KKMS08, GKK08a]. Our work shows that one can achieve an $n^{O(\log n)}$-time result with only quantum examples.
@@
The class of size-$s$ depth-33 circuits is quantum PAC learnable upto error $\varepsilon$ in time $\operatorname{poly}(n,(s/\varepsilon)^{\log\log(s/\varepsilon)\cdot\log(1/\varepsilon)})$.
@@
Agnostic DNF
$n^{\log(1/\varepsilon)\log\log n}$
[GKK08a]
$n^{\log(1/\varepsilon)\log n}$
[KKMS08]
$n^{\log(1/\varepsilon)\log\log n}$
This work
PAC depth-$3$
$n^{\log(1/\varepsilon)\log\log n}$
[Fel09, KK09]
$n^{\log(1/\varepsilon)\log n}$
[Tal17]
$n^{\log(1/\varepsilon)\log\log n}$
This work
@@ 1.2.1 In the quantum setting
Observe that Fourier sampling does work for parity states, denoted as ${\mathcal{C}}_{\mathsf{Par}}=\{\chi_{S}(x)=\langle S,x\rangle\text{ for all }S\in{\{0,1\}}^{n}\}$, since the Fourier spectrum of parity function $\chi_{S}$ is a point-mass on $|S\rangle$. In particular, this implies that if $|\psi\rangle$ is $\tau$-close to $|\chi_{S}\rangle$, then must have a “large” amplitude on $|S\rangle$. Hence, strong learning via Fourier sampling can be achieved using $O(1/\tau^{2})$ samples. But this no longer holds for other classes of states and, in particular, DNFs for which the Fourier spectrum does not enjoy such a “point-concentration” property. One can instead consider Bell sampling, which has been used in [GIKL24a] to agnostically learn stabilizer product states (of which parity states are a subclass). Unfortunately, the results in [GIKL24a] solve a more general problem than the one we are considering and the resulting time and sample complexity of their results is $\operatorname{poly}(n,2^{1/\tau})$. As we are aiming for a polynomial scaling in $1/\tau,$ this is unsatisfactory.
2. Stabilizer Bootstrapping: Another approach is to look at the recent work of Chen et al. [CGYZ25] for agnostic learning of stabilizer states. As with [GIKL24a], however, they solve the more general task of agnostically learning states with stabilizer dimension-$t$, achieving sample and time complexity scaling $\operatorname{poly}(n,(1/\tau)^{\log 1/\tau},2^{t})$. It is not hard to see that learning $s$-juntas or $s$-term DNFs reduces to agnostic learning states with stabilizer dimension $O(s)$, so their agnostic learning algorithm would scale as $\operatorname{poly}(n,(1/\tau)^{\log 1/\tau},2^{s})$.
@@ 1.3.2 Learning algorithms
Agnostic learning DNFs. The agnostic learner for DNFs is similar to the one for decision trees, except that we need to use Mansour’s result [Man92, LT22] that shows that for size-$s$ DNF formulas, the entire Fourier spectrum is concentrated on $s^{O(\log\log s\log(1/\varepsilon))}$ coefficients. Using a similar argument as the one for deriving Eq. (2), one can show that if $|\psi\rangle$ is $\tau$-close to a DNF phase state, then there exists a basis state $|\chi_{T}\rangle$ such that $|\langle\psi|\chi_{T}\rangle|^{2}\geq\tau/s^{O(\log\log s\log(1/\varepsilon))}$. Once again using our quantum boosting algorithm, we obtain a $\operatorname{poly}(n,s^{\log\log s\log(1/\varepsilon)})$ algorithm for agnostic learning size-$s$ DNF formulas.
PAC learning depth-$3$ circuits. To learn these circuits, we employ our agnostic DNF learner. The key insight is that when the input state $|\psi_{f}\rangle$ is promised to correspond to a depth-$3$ circuit then state tomography is accomplished when agnostic learning against DNFs is accomplished, which was also observed classically [Fel09]. This follows from using the seminal result of Hajnal et al. [HMP+93] that says that if $f$ is a threshold of $m$ many DNF formulas $\{g_{1},\ldots,g_{m}\}$ (each with size at most $s$), then $|\langle\psi_{f}|\psi_{g_{i}}\rangle|\geq 1/m,$ for some $i\in[m]$, where $|\psi_{f}\rangle$ (in this section) equals $\frac{1}{\sqrt{2^{n}}}\sumop\displaylimits_{x}f(x)|x\rangle$ since we are in the PAC learning setting. We can now use our agnostic DNF learner outputs a quantum state $|\phi\rangle$ which is at least $\textsf{opt}/m^{2}$ close to $f$. This will serve as our weak learner which we will then boost into a strong PAC learner. The boosting algorithm outputs a (classical description of a) quantum state $|\phi\rangle$ which is close to an unknown $|\psi_{f}\rangle$; we now round the final state $|\phi\rangle$ of the algorithm and show that it satisfies the requirement of PAC learning. Since the runtime of the DNF learner scales as $\operatorname{poly}(n,s^{\log\log s\log(1/\varepsilon)})$, the overall quantum PAC learning algorithm scales as $\operatorname{poly}(n,s^{\log\log s\log(1/\varepsilon)})$.
@@
Finally, we remark that there are two natural definitions of quantum agnostic learning: the one defined in the introduction of this work, i.e., $|\psi\rangle$ is arbitrary and promised to be close to $|\phi_{c}\rangle$ (where $c\in{\mathcal{C}}$), or the one that was considered in [ADW17, CEH+24] wherein there is an unknown distribution $D:{\{0,1\}}^{n+1}\rightarrow[0,1]$ whose first $n$ bits are uniform and the last bit is described by the marginal $\big((1+\phi(x))/2,(1-\phi(x))/2\big)$ where $\phi:{\{0,1\}}^{n}\rightarrow[-1,1]$ is an arbitrary function. The quantum algorithm is given copies of
@@ 1.4 Outlook
Related works. We remark that there have been a few recent works on quantum agnostic learning that we briefly describe here. In [ADW17], the authors showed that quantum examples are equivalent to classical examples in the distribution-independent framework for agnostic learning function classes. Badescu and O’Donnell [BO21] considered the setting of agnostic learning arbitrary classes of quantum states and gave sample complexity bounds for this task. In another direction, works of [CTB24, CHI+23, CEH+24] considered quantum agnostic learning for the case where the input state has some specific structure (in the former work they assume it is a function state and in the latter two works they assume it is a “mixture of superpositions”).
@@
    Learning depth-$3$ circuits: Our PAC learning algorithm for depth-$3$ circuits (with constant $\varepsilon$) runs in “near-polynomial time,” $n^{O(\log\log n)}.$ Is there a $\operatorname{poly}(n)$-time quantum algorithm?
@@
    Learning quantum objects: In this work, We considered the learnability of depth-$2$ and depth-$3$ $\textsf{AC}^{0}$ circuits, what about learning depth-$2$ or depth-$3$ $\textsf{QAC}^{0}$ circuits with or without fanout gates? Recently, [FPVY25] proved the hardness of learning $\textsf{QAC}^{0}$ circuits, but their hard instances require depth that is a “large constant.”
    Similarly, we could consider the problem of agnostic learning low-degree phase states. Tomography protocols for these class of states are known [ABDY23] but agnostic learning algorithms are unknown.
@@
    Similarly, classically the works [JKS02, CLLO21] have considered the class of learning threshold of $\textsf{AC}^{0}$ gates, and [CIKK16] has looked at learning $\textsf{AC}^{0}$ augmented with $\mod p$ gates, with membership queries. It would be interesting if one could learn threshold of $\textsf{AC}^{0}[p]$ circuits using only quantum examples (removing the need for classical queries).
  5. 5.
@@ 2.4.1 PAC learning
In particular, the learner is given copies of $|\psi_{c}\rangle$ and is allowed to perform arbitrary measurements on those copies. Note that measuring $|\psi_{c}\rangle$ in the computational basis produces a classical labeled example, so quantum examples are at least as strong as classical examples. Understanding the strengths and weaknesses of quantum examples has been looked at by several works (we refer an interested reader to the survey [ADW17]). As with the classical complexities, one can similarly define the $(\varepsilon,\delta)$-sample and time complexity for learning $\mathcal{C}$} as the quantum sample complexity (i.e., number of quantum examples $|\psi_{c}\rangle$ used) and quantum time complexity (i.e., number of one and two-qubit quantum gates used in the algorithm) of an optimal $(\varepsilon,\delta)$-learner for $\mathcal{C}$}.

### v3

# Learning depth-3 circuits via quantum agnostic boosting

@@ Abstract
Using quantum agnostic boosting, we give a $n^{O(\log(n/\varepsilon)\cdot\log\log n)}$-time algorithm for $\varepsilon$-learning $\textsf{poly}(n)$-sized depth-$3$ circuits (consisting of AND, OR, NOT gates) in the uniform PAC model given quantum examples. Classically, obtaining an algorithm with a similar complexity has been an open question in the PAC model and our work answers this given quantum examples.
@@ 1 Introduction
Classically, a well-known idea to PAC learn depth-$3$ circuits is to consider the formulation of agnostic learning. In particular, works of [FEL09, KMV08b] showcased how agnostic learning depth-$2$ circuits, in particular DNF formulas, could yield learning algorithms for depth-$3$ circuits. This naturally motivates the need to understand learning DNFs in the quantum agnostic learning model.
@@
Tomography of quantum states (i.e., learning quantum states in the noise-free model) has been well-studied in quantum computing. Quantum agnostic learning got traction only recently with the works of Grewal et al. [GIK+24b] and Chen et al. [CGY+25] that considered learning stabilizer states. Subsequently, a few works considered stabilizer product states [GIK+24a], high stabilizer-dimension states [CGY+25] and product states [BBK+25]. In the pursuit of agnostic learning DNFs, we initiate the problem of agnostic learning phase states (which is incomparable to the works mentioned and which we will discuss in more detail below). We now define the model of agnostic learning specialized to phase states. For the concept class ${\mathcal{C}}\subseteq\{c:{\{0,1\}}^{n}\rightarrow{\{0,1\}}\}$, we denote the phase state $|\phi_{c}\rangle$ corresponding to a function $c\in{\mathcal{C}}$ as
@@ 1.1 Main results
In this work, we make progress on both of the questions highlighted above. For the tasks of agnostically learning $\operatorname{poly}(n)$-sized DNFs, we give a $n^{O(\log\log n\cdot\log(1/\varepsilon))}$-time quantum algorithm and for PAC learning $\operatorname{poly}(n)$-sized depth-$3$ circuits, we give $n^{O(\log(n/\varepsilon)\cdot\log\log n)}$ time quantum algorithm. Below, we will discuss these results, starting with our main technical contribution, quantum agnostic boosting. Classically obtaining similar complexities given only classical examples is an open question and we compare both classical and quantum complexities below.
@@
We formally state this theorem in Section 3. To provide some context, classically, Freund and Schapire [SCH90, FRE95, FS99] proposed the boosting algorithm called AdaBoost that efficiently uses a weak learner as a black-box to construct a strong learner in the usual PAC model. The AdaBoost algorithm by Freund and Schapire was one of the first few theoretical boosting algorithms that was simple enough to be extremely useful and successful in practice [SF12]. Similarly, agnostic boosting has been considered by works of [BLM01, KMV08b, FEL09] wherein they show similar results to boost weak agnostic learners to strong ones.
In the quantum setting, there are only a handful of works that have used boosting in the PAC model [BJ95, AM20, Id20], and ours is the first work that demonstrates how to perform boosting in the harder model of agnostic learning. [^1] We remark that [CTB24] also discusses an agnostic quantum booster but their model assumes that the unknown quantum state is a function state whereas we make no assumption on our input state. Apart from the context of learning Boolean functions, we believe that our quantum boosting algorithm could have utility for learning more general quantum states. Recently, [AD25] utilized boosting on top of a weak agnostic learner for stabilizer states to give tomography protocols of states promised to have structured stabilizer decompositions.
@@
We formally state this theorem for each class in Section 4. In order to compare with classical results and for ease of exposition, let us consider the parameters above to take values of $2^{k},t,s,1/\varepsilon=\operatorname{poly}(n)$ (which in general are considered the most interesting settings of these parameters for learning). Classically, the SOTA algorithm for learning decision trees in the agnostic model (without membership queries) runs in $n^{O(\log n)}$ [KKM+08a] time and with membership queries scales as $\operatorname{poly}(n)$ [GKK08b]. [^3] By membership queries we mean that an algorithm can query an unknown $f:{\{0,1\}}^{n}\rightarrow{\{0,1\}}$ on an $x$ of its choice. In the (uniform) agnostic model, by membership query we mean the following: for every distribution $D:{\{0,1\}}^{n+1}\rightarrow[0,1]$ whose marginal on the first $n$ bits is uniform, the learner can query $x$ and obtain a sample from $b\sim D(x,\cdot)$. The same results hold for juntas since $k$-juntas have size-$2^{k}$ decision trees. In contrast, for DNF formulas, the SOTA algorithm in the agnostic model is $n^{O(\log^{2}n)}$ [KKM+08a, GKK08a]. Our work shows that one can achieve an $n^{O(\log n)}$-time result with only quantum examples.
@@
The class of size-$s$ depth-33 circuits is quantum PAC learnable upto error $\varepsilon$ in time $\operatorname{poly}(n,(s/\varepsilon)^{\log(s/\varepsilon)\cdot\log\log(s/\varepsilon)})$.
@@
Agnostic DNF $n^{\log(1/\varepsilon)\log\log n}$ [GKK08a] $n^{\log(1/\varepsilon)\log n}$ [KKM+08a] $n^{\log(1/\varepsilon)\log\log n}$ This work
PAC depth-$3$ $n^{\log(n/\varepsilon)\log\log n}$ [FEL09, KK09] $n^{\log(n/\varepsilon)\log n}$ [TAL17] $n^{\log(n/\varepsilon)\log\log n}$ This work
@@ 1.2.1 In the quantum setting
Observe that Fourier sampling does work for parity states, denoted as ${\mathcal{C}}_{\mathsf{Par}}=\{\chi_{S}(x)=\langle S,x\rangle\text{ for all }S\in{\{0,1\}}^{n}\}$, since the Fourier spectrum of parity function $\chi_{S}$ is a point-mass on $|S\rangle$. In particular, this implies that if $|\psi\rangle$ is $\tau$-close to $|\chi_{S}\rangle$, then must have a “large” amplitude on $|S\rangle$. Hence, strong learning via Fourier sampling can be achieved using $O(1/\tau^{2})$ samples. But this no longer holds for other classes of states and, in particular, DNFs for which the Fourier spectrum does not enjoy such a “point-concentration” property. One can instead consider Bell sampling, which has been used in [GIK+24a] to agnostically learn stabilizer product states (of which parity states are a subclass). Unfortunately, the results in [GIK+24a] solve a more general problem than the one we are considering and the resulting time and sample complexity of their results is $\operatorname{poly}(n,2^{1/\tau})$. As we are aiming for a polynomial scaling in $1/\tau,$ this is unsatisfactory.
2. Stabilizer Bootstrapping: Another approach is to look at the recent work of Chen et al. [CGY+25] for agnostic learning of stabilizer states. As with [GIK+24a], however, they solve the more general task of agnostically learning states with stabilizer dimension-$t$, achieving sample and time complexity scaling $\operatorname{poly}(n,(1/\tau)^{\log 1/\tau},2^{t})$. It is not hard to see that learning $s$-juntas or $s$-term DNFs reduces to agnostic learning states with stabilizer dimension $O(s)$, so their agnostic learning algorithm would scale as $\operatorname{poly}(n,(1/\tau)^{\log 1/\tau},2^{s})$.
@@ 1.3.2 Learning algorithms
Agnostic learning DNFs. The agnostic learner for DNFs is similar to the one for decision trees, except that we need to use Mansour’s result [MAN92, LT22] that shows that for size-$s$ DNF formulas, the entire Fourier spectrum is concentrated on $s^{O(\log(1/\varepsilon)\cdot\log\log s)}$ coefficients. Using a similar argument as the one for deriving Eq. (2), one can show that if $|\psi\rangle$ is $\tau$-close to a DNF phase state, then there exists a basis state $|\chi_{T}\rangle$ such that $|\langle\psi|\chi_{T}\rangle|^{2}\geq\tau/s^{O(\log(1/\varepsilon)\cdot\log\log s)}$. Once again using our quantum boosting algorithm, we obtain a $\operatorname{poly}(n,s^{\log\log s\log(1/\varepsilon)})$ algorithm for agnostic learning size-$s$ DNF formulas.
PAC learning depth-$3$ circuits. To learn these circuits, we employ our agnostic DNF learner. The key insight is that when the input state $|\psi_{f}\rangle$ is promised to correspond to a depth-$3$ circuit then state tomography is accomplished when agnostic learning against DNFs is accomplished, which was also observed classically [FEL09]. This follows from using the seminal result of Hajnal et al. [HMP+93] that says that if $f$ is a threshold of $m$ many DNF formulas $\{g_{1},\ldots,g_{m}\}$ (each with size at most $s$), then $|\langle\psi_{f}|\psi_{g_{i}}\rangle|\geq 1/m,$ for some $i\in[m]$, where $|\psi_{f}\rangle$ (in this section) equals $\frac{1}{\sqrt{2^{n}}}\sumop\displaylimits_{x}f(x)|x\rangle$ since we are in the PAC learning setting. We can now use our agnostic DNF learner outputs a quantum state $|\phi\rangle$ which is at least $\textsf{opt}/m^{2}$ close to $f$. This will serve as our weak learner which we will then boost into a strong PAC learner. The boosting algorithm outputs a (classical description of a) quantum state $|\phi\rangle$ which is close to an unknown $|\psi_{f}\rangle$; we now round the final state $|\phi\rangle$ of the algorithm and show that it satisfies the requirement of PAC learning. Since the runtime of the DNF learner scales as $\operatorname{poly}(n,s^{\log(1/\varepsilon)\cdot\log\log s})$, the overall $\delta$-error quantum PAC learning algorithm scales as $\operatorname{poly}(n,s^{\log(s/\delta)\cdot\log\log s})$.
@@
Finally, we remark that there are two natural definitions of quantum agnostic learning: the one defined in the introduction of this work, i.e., $|\psi\rangle$ is arbitrary and promised to be close to $|\phi_{c}\rangle$ (where $c\in{\mathcal{C}}$), or the one that was considered in [AD17, CEH+24] wherein there is an unknown distribution $D:{\{0,1\}}^{n+1}\rightarrow[0,1]$ whose first $n$ bits are uniform and the last bit is described by the marginal $\big((1+\phi(x))/2,(1-\phi(x))/2\big)$ where $\phi:{\{0,1\}}^{n}\rightarrow[-1,1]$ is an arbitrary function. The quantum algorithm is given copies of
@@ 1.4 Outlook
Related works. We remark that there have been a few recent works on quantum agnostic learning that we briefly describe here. In [AD17], the authors showed that quantum examples are equivalent to classical examples in the distribution-independent framework for agnostic learning function classes. Badescu and O’Donnell [BO21] considered the setting of agnostic learning arbitrary classes of quantum states and gave sample complexity bounds for this task. In another direction, works of [CTB24, CHI+23, CEH+24] considered quantum agnostic learning for the case where the input state has some specific structure (in the former work they assume it is a function state and in the latter two works they assume it is a “mixture of superpositions”).
@@
    Learning quantum objects: In this work, We considered the learnability of depth-$2$ and depth-$3$ $\textsf{AC}^{0}$ circuits, what about learning depth-$2$ or depth-$3$ $\textsf{QAC}^{0}$ circuits with or without fanout gates? Recently, [FPV+25] proved the hardness of learning $\textsf{QAC}^{0}$ circuits, but their hard instances require depth that is a “large constant.”
    Similarly, we could consider the problem of agnostic learning low-degree phase states. Tomography protocols for these class of states are known [ABD+23] but agnostic learning algorithms are unknown.
@@
    Similarly, classically the works [JKS02, CLL+21] have considered the class of learning threshold of $\textsf{AC}^{0}$ gates, and [CIK+16] has looked at learning $\textsf{AC}^{0}$ augmented with $\mod p$ gates, with membership queries. It would be interesting if one could learn threshold of $\textsf{AC}^{0}[p]$ circuits using only quantum examples (removing the need for classical queries).
@@ 2.4.1 PAC learning
In particular, the learner is given copies of $|\psi_{c}\rangle$ and is allowed to perform arbitrary measurements on those copies. Note that measuring $|\psi_{c}\rangle$ in the computational basis produces a classical labeled example, so quantum examples are at least as strong as classical examples. Understanding the strengths and weaknesses of quantum examples has been looked at by several works (we refer an interested reader to the survey [AD17]). As with the classical complexities, one can similarly define the $(\varepsilon,\delta)$-sample and time complexity for learning $\mathcal{C}$} as the quantum sample complexity (i.e., number of quantum examples $|\psi_{c}\rangle$ used) and quantum time complexity (i.e., number of one and two-qubit quantum gates used in the algorithm) of an optimal $(\varepsilon,\delta)$-learner for $\mathcal{C}$}.